OpenStudy (anonymous):

El Segundo high school put on their annual musical. The students sold 650 tickets for a value of $4375. If orchestra seats cost $7.50 and balcony seats cost $3.50, how many of each kind of seat were sold?

OpenStudy (anonymous):

7.5x+3.5y=4375 x+y=650 y=650-x 7.5x+3.5(650-x)=4375